OpenStudy (anonymous):

Simplify the expression:

OpenStudy (anonymous):

\[144 ^{-1/2}\]

OpenStudy (anonymous):

take the square root take the reciprocal

OpenStudy (maheshmeghwal9):

yeah

OpenStudy (anonymous):

so 12?

OpenStudy (anonymous):

\[144^{-1/2}=(12^2)^{-1/2}\]

OpenStudy (anonymous):

I can only choose from 1/72, -72, 1/12, -12

OpenStudy (anonymous):

1/12, cause (12^2)^(-1/2)=(12)^-1=1/12

OpenStudy (anonymous):

as mukushla said above